OpenStudy (kittykitty04):

Let x=3 tan theta in sqrt (9+ x^2) Find sec theta?

OpenStudy (anonymous):

can you write it down

OpenStudy (kittykitty04):

@Kanwar245

OpenStudy (kittykitty04):

let x = \[3 \tan \theta \] in \[\sqrt{9+x^2}\] find \[\sec \theta \]